OpenStudy (anonymous):

MEDALS!!! Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4-x and y = 2x + 3 intersect are the solutions of the equation 4-x = 2x + 3. Part B: Make tables to find the solution to 4-x = 2x + 3. Take the integer values of x between -3 and 3. Part C: How can you solve the equation 4-x = 2x + 3 graphically?

OpenStudy (anonymous):

Okay check this:

OpenStudy (anonymous):

Part A: To find the point where the two lines intersect, you must find the common point (x). To find this, you must set the two equations equal to each other. I'm not sure how to show you part B

OpenStudy (anonymous):

Ok, makes sense. I believe what B is saying is make tables and it has to range between -3 and 3

OpenStudy (anonymous):

@charlotte123

OpenStudy (charlotte123):

For B Make a table - I shall hand you the Data.

OpenStudy (charlotte123):

4-(1) = 2(1) + 3 4-(2) = 2(2) + 3 4-(3) = 2(3) + 3 4-(0) = 2(0) + 3 4-(-1) = 2(-1) + 3 4-(-2) = 2(-2) + 3 4-(-3) = 2(-3) + 3

OpenStudy (charlotte123):

False - 1st one As 3 ≠ 5 Second one as well - As 2 ≠ 7 3rd one as well - As 1 ≠ 9

OpenStudy (anonymous):

Thank You!!! :):):):):)

OpenStudy (anonymous):

Is that all to Part C???

OpenStudy (anonymous):

Is that all to part C??

OpenStudy (charlotte123):

4th one as well - As 4 ≠ 3 5th one as well - As 5 ≠ 1 6th one as well - As 6 ≠ -1 Last one as well - As 7 ≠ -3

OpenStudy (anonymous):

Is this B or C??

OpenStudy (charlotte123):

For B Hold on

OpenStudy (charlotte123):

Graph for C ----> http://www.wolframalpha.com/input/?i=4-x+%3D+2x+%2B+3+ <----

OpenStudy (anonymous):

Ok thanks but I need to put that in Word

OpenStudy (anonymous):

Or something like that and I can't put graphs like that in Word

OpenStudy (anonymous):

It won't let me!!! UGH!!!!!! Not your fault, but is there any way you can explain that in words?? Please